Cercarial dermatitis, commonly known as swimmers itch, is a rash caused by an allergic reaction to parasites that burrow i
Answer #1

Solution :

n = 142

x = 59

\hat p = x / n = 59 / 142 = 0.415

1 -\hat p = 1 - 0.415 = 0.585

At 98% confidence level the z is ,

\alpha = 1 - 98% = 1 - 0.98 = 0.02

\alpha / 2 = 0.02 / 2 = 0.01

Z\alpha/2 = Z0.01 = 2.326

Margin of error = E = Z\alpha / 2 * \sqrt((\hat p * (1 - \hat p)) / n)

= 2.326 * (\sqrt((0.415 * 0.585) / 142)

= 0.096

A 98 % confidence interval for population proportion p is ,

\hat p- E < P <\hat p + E

0.415 - 0.096 < p < 0.415 + 0.096

0.319 < p < 0.511